OpenStudy (anonymous):

What are the period and phase shift for f(x) = 5 tan(2x − 𝜋)?

hero (hero):

For y = Atan(Bx - C) + D, the period is A and the phase shift is C/B But for the question you posted, the value of C doesn't show up.

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards \\ \quad \\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ % template start y= {\color{purple}{ 5}} ( {\color{blue}{ 2}}x + {\color{red}{ \pi }} ) + {\color{green}{ D}}\\ % template ends \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to\ the\ left \\ \quad \\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to\ the\ right\end{array} \end{array} \\ \quad \\ period\to \cfrac{2\pi }{{\color{blue}{ B}}} }\)